# Jin Li    Ye Luo    Xiaowei Zhang

### v3

@@
Seesaw Experimentation: A/B Tests with Spillovers thanks: Contact information: Jin Li (jli1@hku.hk), Ye Luo (kurtluo@hku.hk), Faculty of Business and Economics, The University of Hong Kong, Pokfulam Road, Hong Kong SAR. Xiaowei Zhang (xiaoweiz@ust.hk), Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong SAR. We thank Shan Huang and Ruohan Zhan for helpful discussions.
@@
(December 31, 2024)
This paper investigates how a firm’s performance can decline even when consistently implementing innovations validated through A/B testing. We introduce the concept of seesaw experimentation, where successful innovations enhance performance in measured primary metrics but generate negative spillover effects in unmeasured secondary dimensions, ultimately reducing overall performance. We identify the conditions under which seesaw experimentation occurs. We also propose a simple solution to address it: implementing a positive hurdle rate for A/B tests. We derive the optimal hurdle rate for both the normal distribution and the fat-tailed distributions.
@@ 1 Introduction
A/B testing has become an important tool for modern business, enabling firms to drive innovation, improve user experiences, and increase revenue growth (Kohavi et al., 2020, Koning et al., 2022). Major tech firms such as Amazon, Google, Meta, and Microsoft deploy thousands of such tests annually, embedding data-driven decision making into their innovation processes (Thomke, 2020, Azevedo et al., 2020).
While existing research typically examines A/B tests in isolation, these experiments frequently generate significant spillover effects [^1] We use spillover, rather than negative spillover, for expositional ease. that ripple across an organization’s ecosystem of products, channels, and performance metrics. These effects are particularly pronounced on large digital platforms. At the WeChat platform, for instance, enhancing Video Accounts can substantially decrease engagement with Official Accounts as users redistribute their limited attention. The spillover effects extends beyond digital interactions: improvements that drive online sales growth often cannibalize offline revenue, and strategies that maximize advertising revenue frequently degrade user experience. These examples underscore a fundamental challenge: improvements in one area of business operations can trigger unintended consequences in seemingly unrelated domains. As externalities are prevalent within organizations, so are the spillover effects.
@@
These spillover effects create what we term seesaw experimentation, where the overall performance of a firm deteriorates despite continuous improvements through innovations in current strategic priorities. The name reflects its oscillating nature. Like a seesaw, advances in one dimension often diminish performance in others. Firms implement innovations that boost their primary metrics—whether in specific products, channels, or KPIs. Yet these same innovations generate hidden or underweighted losses in secondary dimensions. Because business environments and strategic priorities constantly evolve, yesterday’s focus on gross merchandise value may shift to user satisfaction tomorrow. Each shift triggers new waves of innovations that may inadvertently reverse previous gains. The pattern creates a dangerous illusion of progress. Success metrics in current focal areas shine brightly. Meanwhile, the firm’s overall performance quietly degrades.
Managing spillover effects in A/B testing poses a formidable challenge. This is especially true for large organizations running numerous simultaneous experiments. Some firms try to address single-product spillovers through managerial coordination. For instance, user growth teams may challenge revenue-boosting innovations that hurt customer satisfaction. But such efforts often fall short. Several factors drive this coordination failure. First, many potential effects—especially long-term impacts—defy easy quantification. Second, the sheer volume of concurrent experiments makes comprehensive coordination impractical. Third, cross-product and cross-channel effects often remain hidden from view. Despite widespread recognition of their importance for platform health, spillover effects rarely factor into experimental design [^2] Communications with experts at a leading Chinese platform. This creates a persistent gap between the ideal of holistic testing and the reality of fragmented experimentation.
The purpose of this paper is to study A/B testing in the presence of spillover effects, with particular attention to the emergence and implications of seesaw experimentation. We investigate the conditions under which firms, despite achieving continuous improvements in primary performance metrics, may experience deteriorating overall outcomes. Our analysis focuses on two key questions: First, under what conditions does seesaw experimentation emerge as firms improve their primary strategic objectives? Second, how can organizations effectively manage these spillover effects without relying on complex cross-managerial coordination?
@@
We develop a bivariate normal distribution framework to demonstrate the emergence of seesaw experimentation. Our model identifies conditions where firms achieve consistent A/B test successes yet face declining overall performance. Two key factors drive this phenomenon. First, a higher signal-to-noise ratio (the absolute value of mean over standard deviation) increases the likelihood of seesaw experimentation. Second, this effect strengthens as the correlation between dimensions becomes more negative.
@@
Our analysis yields a simple yet powerful solution to manage spillover: implement a positive hurdle rate for A/B test approval. This approach only implements innovations that exceed a specified performance threshold. We derive the optimal hurdle rate where the positive gains in the primary dimension equal the negative externalities in the secondary dimension. Our framework provides practical guidance by showing how this optimal rate varies with key distribution parameters. This solution effectively balances innovation benefits against spillover costs without requiring complex coordination mechanisms.
@@
The literature addressing our specific focus is relatively small. Berman et al., (2018) examine how p-hacking in A/B testing can lead to false discoveries and impair experimental improvements. Azevedo et al., (2020) analyze how distributional characteristics, particularly fat tails, influence experimentation allocation strategies. [^3] We extend our analysis to fat-tailed distributions in Section 4. McClellan, (2022) explores how agency problems affect A/B test adoption mechanisms. While these studies provide valuable insights, they primarily focus on single-dimensional outcomes. Our work departs from this literature by explicitly modeling the interconnectedness of performance across multiple dimensions, revealing how improvements in one dimension can generate spillover effects in others.
@@ 2 Model Formulation
The firm then conducts an A/B test to evaluate a potential innovation using a metric aligned with the primary dimension $a_{t}$, measuring the innovation’s effect in this dimension. Let $u_{t}$ and $v_{t}$ represent the innovation’s effects on dimensions $u$ and $v$, respectively. Note that only the effect on the primary dimension is measured in the A/B test. [^4] This assumption reflects practical challenges: initiators of the A/B test may not have the incentive to measure the outcomes in secondary dimensions, and, in general, outcomes are often harder to measure and may rely on private information held by different stakeholders. But even if firms can perfectly measure outcomes from the secondary dimension, their decisions depend on how to weigh the relative importance of the two. Our analysis reflects the limiting case where the importance of the primary dimension (at the moment of the decision) far outweighs the secondary dimension.
@@
Assumptions 1 and 2 allow us to deliver the main message of the paper—the existence of seesaw experimentation and its mitigation via a positive hurdle rate—in the simplest and most standard framework. The conditions in these assumptions can be relaxed without affecting the main points of the paper. In particular, in Assumption 1(i), $a_{t}$ can be correlated over time, and Assumption 1(ii) is also not necessary in the sense that our main results hold even if the choice of innovation can depend on arbitrary realizations of outcomes in the past. For Assumption 2(i), our results also apply to broader classes of distributions, including those having fat tails and those whose values can on discrete values. Assumption 2(ii) is not necessary in the sense that seesaw experimentation can still occur in scenarios where either $\mu_{u}$ or $\mu_{v}$ is non-negative while the other is negative, provided that $\mu_{u}+\mu_{v}<0$. But we assume this condition to reflect the fundamental business reality that random innovations typically do not improve secondary performance dimensions. Below, we characterize conditions under which seesaw experimentation occurs and analyze how to determine an optimal hurdle rate to address this issue.
3 Base Model
@@
Proposition 1 also provides a sufficient condition for seesaw experimentation to occur. One factor is the signal-to-noise ratio $\alpha=|\mu|/\sigma$. We show in the Appendix (Lemma 1) that $\alpha M(\alpha)$ is an increasing function of $\alpha$, implying that the sufficient condition is more likely to be satisfied when signal-to-noise ratio increases. This can happen in two ways: either through an increase in $|\mu|$ or a decrease in $\sigma$. The intuition behind this is straightforward. When $|\mu|$ is larger, the negative impact on the secondary dimension becomes more pronounced. Alternatively, when $\sigma$ is smaller, there’s less variability in outcomes, which typically results in smaller average benefits in the primary dimension. Both scenarios make seesaw experimentation more likely to occur.
@@
4 Extensions
@@ 4.1 Asymmetry
In addition, the optimal hurdle rate also depends on the mean of the primary dimension, except when the correlation coefficient $\rho$ is zero. In the special case where $\rho=0$, the hurdle rate equals $-\mu$ (the negative externality imposed on the other dimension). When $\rho$ is negative, the optimal hurdle rate $z_{u}^{*}$ decreases as $\mu_{u}$ increases, because a higher mean in the primary dimension reduces the negative externality, leading to a higher optimal hurdle rate. [^5] Under the bivariate normal assumption, the negative externality is $\mathbb{E}[-V|U]=\rho(\sigma_{v}/\sigma_{u})(\mu_{u}-U)-\mu_{v}$. As a result, when $\rho<0$, an increase in $\mu_{u}$ reduces the negative externality. Similarly, when $\rho$ is positive, $z_{u}^{*}$ increases in $\mu_{u}$.
4.2 Multidimensionality
@@
$\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\right]=\sum_{i=1}^{n}\mathbb{P}(A=i)\mathbb{P}(X_{j}>z)\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\bigg{|}X_{j}>z\right]\coloneqq h(z).$ (4)
@@ 4.3 Fat Tails
The parameter $\delta$ controls the tail thickness of the distribution, with smaller values of $\delta$ corresponding to heavier tails. When $\delta=2$, the distribution lacks finite second moments, making covariance and correlation undefined. For $\delta>2$, the scale matrix $\Sigma$ equals the covariance matrix of $(U,V)$, where $\sigma^{2}$ represents the marginal variance and $\rho$ the correlation. As $\delta\to\infty$, the distribution converges to a bivariate normal with the same mean vector and covariance matrix, eliminating the fat tails.
@@
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)W\left(\frac{-\mu}{\sigma}\right)-1,$ (5)
@@
In equation (5), the maximum correlation permitting seesaw experimentation is $12\alpha W(\alpha)-$ 1, where $\alpha=-\mu/\sigma$. Figure 2 illustrates this threshold as a function of $\alpha$ for different values of $\delta$. As $\delta$ increases, this threshold increases for any given $\alpha$. This indicates that, holding marginal mean ($\mu$) and variance ($\sigma^{2}$) constant, lighter tails expand the range of correlations that permit seesaw experimentation, making such phenomenon more likely to occur. This result has an intuitive explanation: with lighter tails, experimentation offers more limited upside potential. Consequently, the negative spillover effects in the secondary dimension are more likely to outweigh the positive gains in the primary dimension.
@@
Hence, the upper bound in equation (5) for the bivariate $t$ distribution converges to the upper bound in (1) for the base model (bivariate normal distribution).
@@
5 Conclusion
Our paper examines A/B testing in the presence of spillover effects. Building on a multivariate normal framework, we identify conditions that give rise to seesaw experimentation—where continuous adoption of successful innovations leads to performance decline. We show that this phenomenon is more likely to occur when the signal-to-noise ratio is large and correlations across dimensions are more negative. We also show how implementing appropriate hurdle rates can enhance firm performance, and derive optimal thresholds within our analytical framework. At the optimal threshold, the positive gain is equal to the negative externality it imposes.
Appendices
Appendix A Proofs for Bivariate Normal Distribution
@@
Let $g(z_{u},z_{v})$ be as defined in (4.1). Then,
$g(z_{u},z_{v})={}$ $p_{u}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right]$ (A.1)
@@
The expression (A.1) can be derived by using the following properties of the truncated normal distribution (Johnson and Kotz, 1972, p. 112–113): for any $z$,
@@
By equation (A.1), a sufficient condition for $g(0,0)<0$ is
$\left\{\begin{array}[]{l}\displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0,\\[10.00002pt] \displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0.\end{array}\right.$ (A.2)
@@
Appendix B Proofs for Multivariate Normal Distribution

### v4

@@
Seesaw Experimentation: A/B Tests with Spillovers thanks: Contact information: Jin Li (jli1@hku.hk), Ye Luo (kurtluo@hku.hk), Faculty of Business and Economics, The University of Hong Kong, Pokfulam Road, Hong Kong SAR. Xiaowei Zhang (xiaoweiz@ust.hk), Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong SAR. We thank Shan Huang, Hanzhe Li, and Ruohan Zhan for helpful comments and discussions.
@@
This paper examines how spillover effects in A/B testing can impede organizational progress and develops strategies for mitigating these challenges. We identify a phenomenon termed seesaw experimentation, where a firm’s overall performance paradoxically deteriorates despite achieving continuous improvements in measured A/B testing metrics. Seesaw experimentation arises when successful innovations in primary metrics generate negative externalities in secondary, unmeasured dimensions. To address this problem, we propose implementing a positive hurdle rate for A/B test approval. We derive the optimal hurdle rate, offering a simple solution that preserves decentralized experimentation while mitigating negative spillovers.
@@ 1 Introduction
A/B testing has become an important tool for modern business, enabling firms to drive innovation, improve user experiences, and increase revenue growth (Kohavi et al., 2020, Koning et al., 2022). Major tech firms such as Amazon, Google, Meta, and Microsoft deploy thousands of such tests annually, embedding data-driven decision-making into their innovation processes (Thomke, 2020, Azevedo et al., 2020).
The massive scale of experimentation in modern organizations requires A/B testing to be initiated with a decentralized structure. While central analytics teams execute these experiments, decisions about what to test and which metrics to track are distributed across product teams, business units, and regional divisions. This decentralization enables rapid innovation but complicates the measurement and management of cross-team effects.
A key challenge of decentralized testing is the occurrence of spillover effects—unintended consequences that arise when teams’ successful A/B test innovations inadvertently harm other parts of the business. These spillovers manifest in various ways: improvements in one product can decrease engagement with others as users redistribute their attention, online sales growth might cannibalize offline revenue, and improving advertising revenue can worsen user experience. Such effects are particularly pronounced on large digital platforms, where products and services are tightly integrated within the same ecosystem. Yet despite their prevalence, these spillover effects remain difficult to measure and often go untracked in the testing process [^1] Industry experts at a leading Chinese digital platform confirm that cross-product externalities are rarely tracked systematically in A/B testing..
The purpose of this paper is to study A/B testing in the presence of spillover effects. We explore how these spillovers [^2] We mostly focus on negative spillover effects in this paper. can fundamentally slow organizational progress and what firms can do to mitigate them. Our first contribution is to highlight and identify seesaw experimentation: a phenomenon in which the overall performance of the firm deteriorates over time even as it achieves continuous improvements in the measured metrics of A/B testing.
@@
The name “seesaw” captures the oscillating nature of this phenomenon. Just as one end of a seesaw rises while the other falls, advances in one dimension often come at the cost of decline in others. When firms implement innovations to boost their primary metrics—whether in specific products, channels, or KPIs—these “improvements” can inflict hidden damage across other areas of the business. As strategic priorities shift, from gross merchandise value today to user satisfaction tomorrow, new innovations may actually reverse previous gains. This creates a dangerous illusion of progress: while the firm is marching from victory to victory, its overall performance quietly declines.
@@
To formally analyze this phenomenon, we develop a bivariate normal distribution framework (later extended to distributions with fat tails) that demonstrates the emergence of seesaw experimentation. Our model identifies specific conditions where firms achieve consistent A/B test successes yet face declining overall performance. Two key factors drive this phenomenon. First, a higher signal-to-noise ratio (the absolute value of mean over standard deviation) increases the likelihood of seesaw experimentation. Second, this effect strengthens as the correlation between dimensions becomes more negative.
@@
Our second contribution is to develop a practical solution for managing spillover effects through a positive hurdle rate for A/B test approval. Rather than implementing all positive innovations, this approach only adopts changes that exceed a specified performance threshold. We analytically derive the optimal hurdle rate that balances gains in the primary dimension against losses in the secondary dimension. The optimal threshold emerges at the point where marginal benefits equal the expected marginal costs in the secondary dimension.
Our framework provides actionable guidance by characterizing how the optimal hurdle rate varies with key parameters of the performance distribution, including the correlation between primary and secondary effects and the means and variance of the distributions. Relative to other potential solutions to deal with the spillover effect, our solution maintains the autonomy of decentralized A/B testing by avoiding the need for complex cross-team coordination mechanisms or comprehensive measurement systems.
@@
A/B testing and experimental design in the digital economy has drawn significant interest from academia and industry; see Zhao, (2024) for a recent survey. However, the literature addressing our specific focus is relatively small. Berman et al., (2018) examine how p-hacking in A/B testing can lead to false discoveries and impair experimental improvements. Azevedo et al., (2020) analyze how distributional characteristics, particularly fat tails, influence experimentation allocation strategies. [^3] We extend our analysis to fat-tailed distributions in the Appendix. McClellan, (2022) explores how agency problems affect A/B test adoption mechanisms. While these studies provide valuable insights, they primarily focus on single-dimensional outcomes. Our work departs from this literature by focusing on the interconnectedness of performance across multiple dimensions, emphasizing how improvements in one dimension can generate spillover effects in others.
Our work also connects to the broader literature on measurement error in economics and organizations. Classical work has long recognized how imperfect measurement can distort decision-making (Heckman, 1979, Holmström and Milgrom, 1991). These measurement challenges arise from both organizational capacity constraints (Ocasio, 1997) and incentive issues (Kerr, 1975). Such challenges become particularly salient in the digital age, where algorithm decision-making can exacerbate the measurement issue (Li et al., 2022, 2023). In our paper, measurement problems at the organizational level can lead the firm to a false sense of progress, even if its performance declines over time.
@@ 2 Model Formulation
The firm then conducts an A/B test to evaluate a potential innovation using a metric aligned with the primary dimension $a_{t}$, measuring the innovation’s effect in this dimension. Let $u_{t}$ and $v_{t}$ represent the innovation’s effects on dimensions $u$ and $v$, respectively. Note that only the effect on the primary dimension is measured in the A/B test.
@@
The model uses several simplifying assumptions to highlight its core mechanism. We now examine how alternative assumptions affect the results.
Remark 1.
Our baseline model assumes one A/B test per period. This assumption can be generalized in several ways. The firm could conduct multiple tests simultaneously across both dimensions, or implement tests probabilistically based on historical outcomes. These extensions preserve our core findings.
Remark 2.
The firm’s adoption rule in our model depends entirely on the primary dimension outcome. When the firm could incorporate both dimensions–such as implementing changes when their weighted average is positive–this would weaken or eliminate the spillover effect. However, the spillover effect persists whenever the firm weighs the primary dimension more heavily, which is likely in practice given business pressures and test initiators’ incentives to prioritize primary metrics.
Remark 3.
We have assumed a bivariate normal distribution, and as we will discuss later, our results extend to settings with multiple dimension and fat-tailed distributions. Within the bivariate normal distribution, Assumptions 1 and 2 can be further relaxed. For example, in Assumption 1, $a_{t}$ can be correlated over time. In addition, Assumption 2(ii) is not essential: seesaw experimentation can still occur when either $\mu_{u}$ or $\mu_{v}$ is non-negative while the other is negative, provided that $\mu_{u}+\mu_{v}<0$.
3 Analysis
@@
Proposition 1 also provides a sufficient condition for seesaw experimentation to occur. One factor is the signal-to-noise ratio $\alpha=|\mu|/\sigma$. (Note: $\mu$ is assumed to be negative.) We show in the Appendix (Lemma 1) that $\alpha M(\alpha)$ is an increasing function of $\alpha$, implying that the sufficient condition is more likely to be satisfied when signal-to-noise ratio increases. This can happen in two ways: either through an increase in $|\mu|$ or a decrease in $\sigma$. The intuition behind this is straightforward. When $|\mu|$ is larger, the negative impact on the secondary dimension becomes more pronounced. Alternatively, when $\sigma$ is smaller, there’s less variability in outcomes, which typically results in smaller average benefits in the primary dimension. Both scenarios make seesaw experimentation more likely to occur.
@@
Remark 4 (Asymmetry).
For asymmetric performance measurements, the firm can apply the same principle as in the symmetric case: set dimension-specific hurdle rates where primary dimension gains exactly offset secondary dimension externalities. The optimal hurdle rate depends on both dimensions’ parameters: notably, with a negative correlation, it decreases as the mean performance in the primary dimension increases, since more gains in one dimension reduce negative externalities in the other.
Remark 5 (Multidimensionality).
Our analysis extends to firms with multidimensional performance measurements. As the number of dimensions increases, seesaw experimentation becomes more prevalent and optimal hurdle rates increase. This follows from the increased likelihood of negative spillovers across dimensions, thereby requiring stricter innovation adoption criteria.
Remark 6 (Fat Tails).
The study of Azevedo et al., (2020) shows that A/B test outcomes often exhibit fat tails that normal distributions fail to capture. We extend our analysis to bivariate $t$ distributions, which generalize normal distributions through a tail-thickness parameter, with lower values producing fatter tails. Under symmetric bivariate $t$ distributions, seesaw experimentation becomes less common with fatter tails, which increase upside potential and thus make negative spillovers less likely to outweigh gains in the primary dimension. The optimal hurdle rate, however, remains unchanged from the normal model. This is because both models yield identical expected externality expressions, and the optimal hurdle rate balances gains in the primary dimension against negative externalities in the secondary dimension (Figure 1).
4 Conclusion
Our paper examines A/B testing in the presence of spillover effects. Building on a bivariate normal framework, we identify conditions that give rise to seesaw experimentation—where continuous adoption of successful innovations leads to performance decline. We show that this phenomenon is more likely to occur when the signal-to-noise ratio is large and correlations across dimensions are more negative. We also show how implementing appropriate hurdle rates can enhance firm performance, and derive optimal thresholds within our analytical framework. At the optimal threshold, the positive gain is equal to the negative externality it imposes.
Appendices
Appendix A Extensions
@@ A.1 Asymmetry
In addition, the optimal hurdle rate also depends on the mean of the primary dimension, except when the correlation coefficient $\rho$ equals zero. In the special case where $\rho=0$, the optimal hurdle rate in one dimension exactly offsets the negative externality imposed on the other (e.g., $z_{u}^{*}=-\mu_{v}$). When $\rho$ is negative, the optimal hurdle rate $z_{u}^{*}$ decreases as $\mu_{u}$ increases, because a higher mean in the primary dimension reduces the negative externality, leading to a higher optimal hurdle rate. [^4] Under the bivariate normal assumption, the negative externality is $\mathbb{E}[-V|U]=\rho(\sigma_{v}/\sigma_{u})(\mu_{u}-U)-\mu_{v}$. As a result, when $\rho<0$, an increase in $\mu_{u}$ reduces the negative externality. Similarly, when $\rho$ is positive, $z_{u}^{*}$ increases in $\mu_{u}$.
A.2 Multidimensionality
@@
$\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\right]=\sum_{i=1}^{n}\mathbb{P}(A=i)\mathbb{P}(X_{j}>z)\mathbb{E}\left[D\sum_{i=1}^{n}X_{i}\bigg{|}X_{j}>z\right]\coloneqq h(z).$ (A.2)
@@ A.3 Fat Tails
The parameter $\delta$ controls the tail thickness of the distribution, with smaller values of $\delta$ corresponding to heavier tails. When $\delta\leq 2$, the distribution lacks finite second moments, making covariance and correlation undefined. For $\delta>2$, the scale matrix $\Sigma$ equals the covariance matrix of $(U,V)$, where $\sigma^{2}$ represents the marginal variance and $\rho$ the correlation. As $\delta\to\infty$, the distribution converges to a bivariate normal with the same mean vector and covariance matrix, eliminating the fat tails.
@@
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)W\left(\frac{-\mu}{\sigma}\right)-1,$ (A.3)
@@
In equation (A.3), the maximum correlation that permits seesaw experimentation is $12\alpha W(\alpha)-$ 1, where $\alpha=-\mu/\sigma$. Figure 2 illustrates this threshold as a function of $\alpha$ for different values of $\delta$. As $\delta$ increases, this threshold increases for any given $\alpha$. This indicates that, holding marginal mean ($\mu$) and variance ($\sigma^{2}$) constant, lighter tails expand the range of correlations that permit seesaw experimentation, making such phenomenon more likely to occur. This result has an intuitive explanation: with lighter tails, experimentation offers more limited upside potential. Consequently, the negative spillover effects in the secondary dimension are more likely to outweigh the positive gains in the primary dimension.
@@
Hence, the upper bound in equation (A.3) for the bivariate $t$ distribution converges to the upper bound in (1) for the base model (bivariate normal distribution).
@@
Appendix B Technical Proofs
B.1 Proofs for Bivariate Normal Distribution
@@
Let $g(z_{u},z_{v})$ be as defined in (A.1). Then,
$g(z_{u},z_{v})={}$ $p_{u}\left[(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{z_{u}-\mu_{u}}{\sigma_{u}}\right)\right]$ (B.1)
@@
The expression (B.1) can be derived by using the following properties of the truncated normal distribution (Johnson and Kotz, 1972, p. 112–113): for any $z$,
@@
By equation (B.1), a sufficient condition for $g(0,0)<0$ is
$\left\{\begin{array}[]{l}\displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)\right)+\left(\sigma_{u}+\rho\sigma_{v}\right)\phi\left(\frac{-\mu_{u}}{\sigma_{u}}\right)<0,\\[10.00002pt] \displaystyle(\mu_{u}+\mu_{v})\left(1-\Phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)\right)+\left(\rho\sigma_{u}+\sigma_{v}\right)\phi\left(\frac{-\mu_{v}}{\sigma_{v}}\right)<0.\end{array}\right.$ (B.2)
@@
B.2 Proofs for Multivariate Normal Distribution